OpenStudy (anonymous):

find the x-intercept for the equation f(x)= 3(x-6)^2+3

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

can anyone solve this quickly

OpenStudy (anonymous):

To find x-intercepts, let y = 0 so, 3(x-6)^2 + 3 = 0 Subtract 3 from both sides, you get 3(x-6)^2 = -3 Divide both sides by 3, you get (x-6)^2 = -1 But no number squared is equal to -1, therefore, there is no real solution to the equation, and hence, no x-intercepts.